OpenStudy (anonymous):

4p^2+12pq+9q^2 factor completely

OpenStudy (anonymous):

\[(2 p+3 q)^2 \]

OpenStudy (anonymous):

? how'd you get that? I got something completely different

OpenStudy (anonymous):

Simplifying 4p2 + 12pq + 9q2 = 0 Reorder the terms: 12pq + 4p2 + 9q2 = 0 Solving 12pq + 4p2 + 9q2 = 0 Solving for variable 'p'. Factor a trinomial. (2p + 3q)(2p + 3q) = 0 Subproblem 1Set the factor '(2p + 3q)' equal to zero and attempt to solve: Simplifying 2p + 3q = 0 Solving 2p + 3q = 0 Move all terms containing p to the left, all other terms to the right. Add '-3q' to each side of the equation. 2p + 3q + -3q = 0 + -3q Combine like terms: 3q + -3q = 0 2p + 0 = 0 + -3q 2p = 0 + -3q Remove the zero: 2p = -3q Divide each side by '2'. p = -1.5q Simplifying p = -1.5q Subproblem 2Set the factor '(2p + 3q)' equal to zero and attempt to solve: Simplifying 2p + 3q = 0 Solving 2p + 3q = 0 Move all terms containing p to the left, all other terms to the right. Add '-3q' to each side of the equation. 2p + 3q + -3q = 0 + -3q Combine like terms: 3q + -3q = 0 2p + 0 = 0 + -3q 2p = 0 + -3q Remove the zero: 2p = -3q Divide each side by '2'. p = -1.5q Simplifying p = -1.5qSolutionp = {-1.5q, -1.5q}